OpenStudy (anonymous):

Factor b12 + 8. (b4- 2)(b16+ 2b4+ 4) (b4+ 2)(b8- 2b4+ 4) (b4- 2)(b8+ 2b4+ 4) (b4+ 2)(b16- 2b4+ 4)

OpenStudy (anonymous):

it is b^12 +8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can write it as: \[(b^4)^3 + (2)^3\] Now use the formula: \[x^3 + y^3 = (x+y)(x^2 - xy + y^2)\] Here: \[x = b^4 \quad and \quad y = 2\]